# Extended State Observer for Localized Fault Awareness in RF Accelerating Structures

### 01 Apr 2026 v1

@@ I Introduction
Radio-frequency (RF) accelerating cavities are fundamental components for transferring energy in modern particle accelerators. These resonant electromagnetic structures store RF energy synchronized with the arrival of charged particle bunches to efficiently transfer energy from high-power RF amplifiers such as diacrodes, klystrons, or solid-state transmitters to the beam [55, 26, 6]. Chains of coupled resonant cavities form accelerating structures through which the electromagnetic behavior is accurately modeled using the theory of coupled resonators [24, 10]. These models describe dispersion characteristics, field stability, and sensitivity to fabrication errors in multicell cavity structures. Because the energy gain imparted to the beam depends directly on the amplitude and phase of the accelerating field, maintaining stable cavity fields is essential for preserving quality beam and minimizing energy spread during accelerator operation [41, 3].
The Los Alamos Neutron Science Center (LANSCE) provides a prominent example of an RF linear accelerator (linac) operating under such conditions. The LANSCE linac accelerates protons to energies approaching 800 MeV using a sequence of RF accelerating structures operating at multiple frequencies [49, 21]. Acceleration begins with a 201.25 MHz drift-tube linac (DTL) that raises the beam energy to approximately 100 MeV [58]. This is followed by a series of 805 MHz side-coupled cavity modules that provide the remaining energy gain [13]. The side-coupled architecture enables stable $\pi/2$-mode operation in long cavity chains while maintaining efficient RF coupling and field uniformity across the structure [50, 24, 13]. Over several decades of operation, LANSCE has provided extensive operational experience with the long-term behavior of RF systems in demanding accelerator environments [48, 25]. As illustrated in Figure 1, LANSCE operates in pulsed mode at 120 Hz with each RF macropulse approximately 1 ms in duration [28].
@@
A distinctive feature of LANSCE operation is the wide diversity of beam flavors delivered to multiple experimental facilities, including the Lujan Neutron Scattering Center, the Proton Radiography Facility, the Weapons Neutron Research Facility, the Isotope Production Facility, and the Ultra Cold Neutron Facility [36, 51]. These programs require beam pulses with substantially different average currents, repetition rates, and time structures [1]. As a result, the RF cavities experience a wide range of beam-loading conditions and disturbances in cavity voltage and phase that depend on beam current and its phase relative to the RF waveform [14]. Additional perturbations arise from cavity detuning caused by thermal expansion, Lorentz-force deformation, and mechanical vibrations. These shift the resonant frequency and modify the efficiency of RF power transfer to the cavity fields [6, 46]. In addition to intrinsic cavity perturbations, temperature variations in transmission cables and electronic components can shift the phases of RF drive and pickup signals relative to the reference. Ongoing modernization efforts at LANSCE continue to upgrade RF infrastructure and digital low-level RF (LLRF) control systems to improve stability and meet the demands of an increasingly wide range of beam conditions [2, 15, 33, 53, 20].
@@
Early accelerator facilities employed analog feedback loops and feedforward compensation to mitigate beam-loading transients and maintain acceptable field stability in pulsed linacs. Advances in digital signal processing and field-programmable gate array (FPGA) platforms have since enabled more flexible and precise digital LLRF architectures [8, 17, 11, 56]. Modern systems represent cavity signals in complex baseband (IQ) form derived by approximate RLC circuit models to enable vector control of cavity fields using high-speed digital feedback loops [22, 39, 18]. These architectures integrate RF detection, digital control algorithms, and timing synchronization to achieve amplitude and phase stability with precision for demanding accelerator environments [42, 43]. Large accelerator facilities such as the TESLA linear collider, the European XFEL, and the LCLS-II free-electron laser have further driven advances in LLRF technology, including vector-sum control of multiple cavities, high-precision RF reference distribution systems, and advanced RF diagnostics [5, 38, 56, 9, 44, 45].
@@
Proportional–integral (PI) control combined with feedforward compensation is widely employed in digital LLRF systems as the basis for feedback regulation of cavity field amplitude and phase [42, 4, 43, 18, 15]. Optimal control has been investigated to incorporate cavity dynamics more directly into the control design for improved transient behavior [38, 40]. At the same time, adaptive and model-independent control methods have been explored to address time-varying operating conditions and uncertainties in RF cavity systems. Techniques based on extremum seeking have been applied to automatically tune cavity resonance and optimize RF system performance without requiring detailed models of the underlying dynamics [34, 32]. Adaptive resonance control and beam-loading compensation strategies have also been proposed for operation under changing beam conditions [33, 36]. These developments reflect a broader trend toward RF control systems capable of accommodating multiple sources of variation while providing improved insight into system behavior [35, 37].
@@
A persistent challenge in RF cavity control is distinguishing between physical sources of phase and amplitude error that produce similar signatures in measured RF signals. In digital LLRF systems, cavity detuning and phase drifts in the forward drive and receiver chains can all appear as rotations of the cavity phasor in the complex baseband representation [19, 57, 12, 31, 7]. Conventional feedback controllers regulate the cavity field without explicitly identifying the origin of these disturbances, which complicates diagnostics and fault detection during accelerator operation. This work develops an observer framework that separates these disturbance mechanisms within a unified baseband cavity model by augmenting the standard cavity field observer with additional states that estimate forward-chain phase drift, cavity detuning, and receiver-chain phase drift. Beyond cavity field regulation, this approach provides a means to identify both the source and location of disturbances within the RF system for more informed diagnostics and troubleshooting before performance degrades or protective interlocks are triggered.
@@ II Baseband Cavity Model
where $V^{\mathrm{cav}}(t)$ is a slowly varying complex envelope whose magnitude and phase determine the accelerating gradient and RF phase relative to the LLRF reference [23]. This section derives a complex baseband state-space model for estimating $V^{\mathrm{cav}}(t)$. The objective is to separate slow drifts introduced by the forward drive and receiver measurement chains from detuning dynamics associated with cavity operation off resonance. Figure 2 illustrates the LLRF feedback loop and channels through which disturbances enter the system.
@@ II-A Drive Synthesis and Forward Path
In digital LLRF systems such as those at LANSCE, a complex baseband voltage command $V(t)$ is synthesized by an FPGA within a field control module (FCM) or equivalent hardware. This signal is converted by a DAC to an analog intermediate frequency (IF), upconverted to the RF carrier, amplified by the high-power amplifier chain, and delivered to the cavity input coupler by waveguide or coaxial transmission. Temperature variations and slow phase drift in the modulator, upconverter, amplifiers, and transmission lines introduce a phase rotation between the commanded signal and the RF drive that reaches the cavity. We model this hardware drift in the forward drive chain as a lumped phase rotation $\varphi^{\mathrm{fwd}}(t)$ between the commanded drive $V(t)$ and the forward voltage $V^{\mathrm{fwd}}(t)$ incident on the cavity according to
@@ II-B Cavity Voltage and Detuning
Detuning relative to the reference is defined by
@@
where $\omega_{r}(t)$ is the instantaneous cavity resonance frequency. When the cavity is on resonance, reflected power from the cavity is minimized for the given coupling. As detuning increases in magnitude, the cavity impedance becomes increasingly reactive. This results in higher reflected power and greater forward RF power required to maintain a specified cavity voltage. Therefore, sustained operation off resonance increases RF power demand and can quickly approach the limits of the high-power amplifiers. At LANSCE, measurements of reflected power are routed to the resonance control (RC) system and the RF interface control (RFIC) chassis for monitoring. The RC system adjusts the cavity frequency by mechanically tuning slugs to minimize reflected power and maintain operation near resonance.
@@
where $\omega_{1/2}=\omega_{0}/2Q_{L}$ is the loaded half-bandwidth and $d(t)$ is a complex disturbance normalized consistently with the voltage definition. This disturbance represents the aggregate effects of beam loading, calibration error, and residual mismatch. The reflected waveform is digitized and digitally downconverted in the FCM to produce complex IQ measurements consistent with the cavity pickup and forward signals. These measurements may be used to establish a systematic method to estimate detuning and distinguish it from hardware phase drifts accumulated along the drive and receiver chains. Incorporating the estimate into the control architecture not only improves drive prediction but increases operator awareness of developing faults that manifest as increasing reflected power.
@@ II-C Pickup and Receiver Chain
The cavity field is monitored by a weakly coupled pickup probe inside the cavity with which a signal proportional to the internal accelerating field is measured. The pickup analog RF signal is downconverted to intermediate frequency (IF) near the cavity to reduce transmission loss and phase drift prior to transport to the FCM. The IF signal is filtered and digitally downconverted to baseband in the FPGA to produce complex IQ signals that signify the amplitude and phase of the cavity relative to the master reference. These IQ signals are used to compute the feedback drive command that regulates the cavity field and closes the control loop.
@@ II-D State-Space Formulation
The IQ planar rotation matrix of sufficiently small angle $\theta(t)$ is approximated by
@@
where the state, control, and measurement matrices are given by
@@
where indices represent the sequence obtained by discrete time sampling. Let us note that the discrete matrices depend on the sampling frequency in the usual way of discretization. In operation at LANSCE, the commanded input $u_{k}$ is held constant over each time segment $[k,k+1)$ of the feedback sequence. Although matrix exponentials may be derived in closed form to represent the exact evolution under piecewise constant control, we assume that the sampling frequency is sufficiently fast to accurately model dynamics with first-order sampling methods. Besides, any residual mismatch is lumped into the additive disturbance.
@@ III-B Forward Drift Observer
where $\alpha_{\mathrm{fwd}}\geq 0$ is the observer gain. The estimated forward drift is applied to the outgoing drive command through the updated control matrix estimate so that the cavity input remains aligned with the controller reference despite hardware phase rotation.
@@ IV-A Feedforward
To drive the state estimation to the setpoint in one step, substitute $\hat{x}_{k+1}=x_{k+1}^{*}$ into (10) and invert $\hat{B}_{k}$ to obtain the feedforward command [16, 52]
@@ V-A Monte Carlo Construction
Each MC realization simulates the discrete cavity model and control architecture over a macropulse containing a fill interval followed by a regulated flattop that collectively span 1 ms. The simulations use ten thousand MC realizations to ensure sufficient exposure of systematic differences between the observer architectures. System parameters are documented in Tables I and II. These are fixed across all MC trials for both the proposed and standard observers. For each MC trial, new randomized truth profiles are generated for the drift channels and noise processes. Detuning is generated using a composite process consisting of a bias term, several sinusoidal components with randomized frequency and amplitude, low-frequency stochastic wander, and a thermal contribution derived from mixer temperature variations. Phase drifts in the forward and receiver chains are generated by slowly evolving stochastic trajectories with small periodic components to represent variations in the RF drive and receiver chains. Trajectories of the additive disturbance vector in the baseband model are generated by similar random processes. Gaussian measurement noise is injected into the cavity pickup, reflected-wave measurement, reference, and forward-drive channels each with a standard deviation of 1e-4. Ten of the ten thousand MC drifts and detuning trajectories as well as their estimations using both the proposed and standard observers are shown in Figure 3.
@@ V-B Regulation
Figure 4: Likelihood that average amplitude and phase regulation errors over time exceed specified thresholds.
@@ V-C Localized Fault Identification
Figure 5: Likelihood that average error over time between true and estimated disturbances exceeds specified thresholds.
@@ VI Conclusion
Several directions for future work remain. Experimental validation of the proposed observer on operational RF stations would provide valuable insight into the interaction between estimator dynamics and real hardware nonlinearities. Additional extensions could incorporate more detailed mechanical models of cavity detuning, including microphonic vibration modes and Lorentz-force effects, which are particularly relevant in superconducting accelerators [19, 47, 57]. Stability of the feedback control should be rigorously established, and adaptive observer gains may be considered for operations that require even tighter stability margins. Finally, integration of the observer outputs into higher-level machine protection and supervisory control systems may enable automated detection of developing RF faults before they impact accelerator performance. As accelerator facilities continue to operate at higher beam power and tighter field stability requirements, the ability to regulate RF cavities while simultaneously diagnosing the origin of disturbances will become increasingly important. Observer-based approaches that explicitly model the underlying physics of RF systems provide a promising path toward achieving both goals.

### 04 Apr 2026 v2

@@ I Introduction
Radio-frequency (RF) accelerating cavities are fundamental components for transferring energy in modern particle accelerators. These resonant electromagnetic structures store RF energy synchronized with the arrival of charged particle bunches to efficiently transfer energy from high-power RF amplifiers such as klystrons and solid-state transmitters to the beam [47, 17]. Chains of coupled resonant cavities form accelerating structures through which the electromagnetic behavior is accurately modeled using the theory of coupled resonators [15, 16]. These models describe dispersion characteristics, field stability, and sensitivity to fabrication errors in multicell cavity structures. Because the energy gain imparted to the beam depends directly on the amplitude and phase of the accelerating field, maintaining stable cavity fields is essential for preserving quality beam and minimizing energy spread during accelerator operation [38].
The Los Alamos Neutron Science Center (LANSCE) provides a prominent example of an RF linear accelerator (linac) operating under such conditions. The LANSCE linac accelerates protons to energies approaching 800 MeV using a sequence of RF accelerating structures operating at multiple frequencies [41, 51]. Following Cockcroft Walton injectors and particle bunchers, acceleration begins with a 201.25 MHz drift-tube linac (DTL) that raises beam energy to approximately 100 MeV [51]. This is followed by a series of 805 MHz side-coupled cavity modules that provide the remaining energy gain [39]. The side-coupled architecture enables stable $\pi/2$-mode operation in long cavity chains while maintaining efficient RF coupling and field uniformity across the structure [39, 40]. Over several decades of operation, LANSCE has provided extensive experience with the long-term operational behavior of RF systems in demanding accelerator environments [7, 13]. As illustrated in Figure 1, LANSCE operates in pulsed mode up to 120 Hz, where RF power is enabled only during the fill and flattop portions of each macropulse.
@@
A distinctive feature of LANSCE operation is the wide diversity of beam flavors delivered to multiple experimental facilities, including the Lujan Neutron Scattering Center, the Proton Radiography Facility, the Weapons Neutron Research Facility, the Isotope Production Facility, and the Ultra Cold Neutron Facility [31, 42]. These programs require beam pulses with substantially different average currents, repetition rates, and time structures [1]. As a result, the RF cavities experience a wide range of beam-loading conditions and disturbances that depend on beam current and its phase relative to the RF waveform [2, 27]. Additional perturbations arise from cavity detuning caused by thermal expansion, Lorentz-force deformation, and mechanical vibrations. These shift the resonant frequency and modify the efficiency of RF power transfer to the cavity fields [11]. In addition to intrinsic cavity perturbations, temperature variations in transmission cables and electronic components can shift the phases of RF drive and pickup signals relative to the reference. Projects at LANSCE continue to upgrade RF infrastructure and digital low-level RF (LLRF) control systems to improve stability and meet the demands of an increasingly wide range of beam flavors [24, 21, 43, 14].
@@
Early accelerator facilities employed analog feedback loops and feedforward compensation to mitigate beam-loading transients and maintain acceptable field stability in pulsed linacs. Advances in digital signal processing and field-programmable gate array (FPGA) platforms have since enabled more flexible and precise digital LLRF architectures [5, 9, 48]. Modern systems represent field signals in complex baseband (IQ) form derived by approximate RLC circuit models to enable vector control of cavity fields using high-speed digital feedback loops [33, 28, 10, 6]. These architectures integrate RF detection, digital control algorithms, and timing synchronization to achieve amplitude and phase stability with precision for demanding accelerator environments. Large accelerator projects such as the International Linear Collider and the European X-Ray Free Electron Laser (XFEL) have more recently provided opportunity to drive advances in RF technology for high-precision RF reference distribution, advanced RF diagnostics, and energy upgrades beyond 1 to 3 TeV [34, 37, 45, 35].
@@
Proportional–integral (PI) control combined with feedforward compensation is widely used in digital LLRF systems as the basis for feedback regulation of cavity field amplitude and phase [3, 38, 10]. Robust and model-predictive control techniques have been proposed to reject disturbances and incorporate cavity dynamics more directly into the control design [19, 49, 46]. At the same time, adaptive and model-independent control methods have been explored to address time-varying operating conditions and uncertainties in RF cavity systems. Techniques based on extremum seeking have been applied to automatically tune cavity resonance and optimize RF system performance without requiring detailed models of the underlying dynamics [30, 29]. Adaptive resonance control and beam-loading compensation strategies have also been proposed for operation under changing beam conditions [31, 27]. These developments reflect a broader trend toward RF control systems capable of accommodating multiple sources of variation while providing improved insight into system behavior [32, 38].
@@
A persistent challenge in RF cavity control is distinguishing between physical sources of phase and amplitude error that produce similar signatures in measured RF signals. In digital LLRF systems, cavity detuning and phase drifts in the forward drive and receiver chains can all appear as rotations of the cavity phasor in the complex baseband representation [12, 50, 25, 4]. Conventional feedback controllers regulate the cavity field without explicitly identifying the origin of these disturbances, which complicates diagnostics and fault detection during accelerator operation. This work develops an observer framework that separates these disturbance mechanisms within a unified baseband cavity model by augmenting the standard cavity field observer with additional states that estimate forward-chain phase drift, cavity detuning, and receiver-chain phase drift. Beyond cavity field regulation, this approach provides a means to identify both the source and location of disturbances within the RF system for more informed diagnostics and troubleshooting before performance degrades or protective interlocks are triggered.
@@ II Baseband Cavity Model
where $V^{\mathrm{cav}}(t)$ is a slowly varying complex envelope whose magnitude and phase determine the accelerating gradient and RF phase relative to the LLRF reference. This section derives a complex baseband state-space model for estimating $V^{\mathrm{cav}}(t)$. The objective is to separate slow drifts introduced by the forward drive and receiver measurement chains from detuning dynamics associated with cavity operation off resonance. Figure 2 illustrates the LLRF feedback loop and channels through which disturbances enter the system.
@@ II-A Drive Synthesis and Forward Path
In digital LLRF systems such as those at LANSCE, a complex baseband voltage command $V(t)$ is synthesized by an FPGA within a field control module (FCM) or equivalent hardware. This signal is converted by a DAC (digital-to-analog converter) to an analog intermediate frequency (IF), upconverted to the RF carrier, amplified by the high-power amplifier chain, and delivered to the cavity input coupler by waveguide or coaxial transmission. Temperature variations and slow phase drift in the modulator, upconverter, amplifiers, and transmission lines introduce a phase rotation between the commanded signal and the RF drive that reaches the cavity. We model this hardware drift in the forward drive chain as a lumped phase rotation $\varphi^{\mathrm{fwd}}(t)$ between the commanded drive $V(t)$ and the forward voltage $V^{\mathrm{fwd}}(t)$ incident on the cavity according to
@@ II-B Cavity Voltage and Detuning
Detuning plays a critical role in RF cavity performance because even small frequency offsets from resonance can significantly impact power efficiency and system stability. Detuning relative to the reference is defined by
@@
where $\omega_{r}(t)$ is the instantaneous cavity resonance frequency. When the cavity is on resonance, reflected power is minimized for the given coupling. As the magnitude of detuning increases, the cavity impedance becomes increasingly reactive, resulting in higher reflected power and greater forward RF power required to maintain a specified cavity voltage. Consequently, sustained operation off resonance can quickly compromise high-power amplifier limits. At LANSCE, measurements of reflected power are routed to the resonance control (RC) system and the RF interface control (RFIC) chassis for monitoring. The RC system mechanically drives tuning slugs to minimize reflected power and maintain operation near resonance.
@@
where $\omega_{1/2}=\omega_{0}/2Q_{L}$ is the loaded half-bandwidth and $d(t)$ is a complex disturbance normalized consistently with the voltage definition. This disturbance represents the aggregate effects of beam loading [26], calibration error, and residual mismatch. The reflected waveform is digitized and digitally downconverted in the FCM to produce complex IQ measurements consistent with the cavity pickup and forward signals. These measurements may be used to establish a systematic method to estimate detuning and distinguish it from hardware phase drifts accumulated along the drive and receiver chains. Incorporating the estimate into the control architecture not only improves drive prediction but increases operator awareness of developing faults that manifest as increasing reflected power.
@@ II-C Pickup and Receiver Chain
The cavity field is monitored by a weakly coupled pickup probe inside the cavity with which a signal proportional to the internal accelerating field is measured. The pickup analog RF signal is downconverted to IF near the cavity input coupler to reduce transmission loss and phase drift prior to transport to the FCM. The IF analog signal is converted by an ADC (analog-to-digital converter) to digital IF, filtered, and digitally downconverted to baseband in the FCM to produce complex IQ phasor signals that signify the amplitude and phase of the cavity field relative to the master reference. These IQ signals are supplied to the FPGA to compute the feedback drive command that regulates the cavity field and closes the control loop.
@@ II-D State-Space Formulation
The IQ planar rotation matrix of a sufficiently small angle $\theta(t)$ is approximated by
@@
where state, control, and measurement matrices are given by
@@
where indices represent the discrete time samples. In operation at LANSCE, the commanded input $u_{k}$ is held constant over each time segment $[k,k+1)$ of the feedback sequence. Although matrix exponentials may be derived in closed form to represent the exact evolution under piecewise constant control, we assume that the sampling frequency is sufficiently fast to accurately model dynamics with first-order sampling methods. Besides, any residual mismatch is lumped into the additive disturbance.
@@ III-B Forward Drift Observer
where $\alpha_{\mathrm{fwd}}\geq 0$ is the observer gain. The estimated forward drift is applied to the drive command through the control matrix update so that the cavity input remains aligned with the controller reference despite hardware phase rotation.
@@ IV-A Feedforward
To drive the state estimation to the setpoint in one step, substitute $\hat{x}_{k+1}^{\mathrm{pred}}=x_{k+1}^{*}$ into (10) and invert $\hat{B}_{k}$ to obtain the feedforward command [8]
@@ V-A Monte Carlo Construction
Each MC realization simulates the discrete cavity model and control architecture over the fill and flattop of a macropulse spanning 1 ms. The simulations use ten thousand MC realizations to ensure sufficient exposure of systematic differences between the observer architectures. System parameters are documented in Tables I and II. These are fixed across all MC trials for both the proposed and standard observers. For each MC trial, new randomized truth profiles are generated for the drift channels and noise processes. Detuning is generated using a composite process consisting of a bias term, several sinusoidal components with randomized frequency and amplitude, low-frequency stochastic wander, and a thermal contribution derived from mixer temperature variations. Phase drifts in the forward and receiver chains are generated by slowly evolving stochastic trajectories with small periodic components to represent variations in the RF drive and receiver chains. Trajectories of the additive disturbance vector in the baseband model are generated by similar random processes. Gaussian measurement noise is injected into the cavity pickup, reflected-wave measurement, reference, and forward-drive channels each with a standard deviation of 1e-4. Ten of the ten thousand MC drifts and detuning trajectories as well as their estimations using both the proposed and standard observers are shown in Figure 3.
@@ V-B Regulation
Figure 4: Likelihood that the average amplitude and phase regulation errors over time exceed the thresholds specified.
@@ V-C Localized Fault Identification
Figure 5: Likelihood that the average error over time between the true and estimated disturbances exceeds the specified thresholds.
@@ VI Conclusion
Several directions for future work remain. Experimental validation of the proposed observer on operational RF stations would provide valuable insight into the interaction between estimator dynamics and real hardware nonlinearities. Additional extensions could consider mechanical models of cavity detuning, including microphonic vibration modes and Lorentz-force effects, which are particularly relevant in superconducting accelerators [18, 12, 36, 50]. Stability of the feedback control should be rigorously established, and adaptive observer gains may be considered for operations that require even tighter stability margins. Finally, integration of the observer outputs into higher-level machine protection and supervisory control systems may enable automated detection of developing RF faults before they impact accelerator performance. As accelerator facilities continue to operate at higher beam power and tighter field stability requirements, the ability to regulate RF cavities while simultaneously diagnosing the origin of disturbances will become increasingly important. Observer-based approaches that explicitly model the underlying physics of RF systems provide a promising path toward achieving both goals.